# Uniqueness of asymptotically cylindrical steady gradient Ricci solitons

### v1

@@ 1.6. Outline of the paper
In §2, we collect relevant facts (due to other authors) about asymptotically $\mathcal{S}$-cylindrical steady GRSs. In particular, any such soliton $(M,g,f)$ opens up like a paraboloid, and the level sets of $f$ converge smoothly (after a certain rescaling) to a spherical space form $(\mathcal{S},\bar{g}_{\infty})$.
@@
where $\operatorname{Rm}(h)_{ik}=\operatorname{Rm}_{ijkl}h^{jl}$ is the action of the curvature tensor. Brendle and Zhao–Zhu [zz] carried this out under extra curvature hypotheses. In contrast, our Liouville theorem (Proposition 3.6) is specific to the tensor $h=\mathcal{L}_{W}g$ but requires no curvature hypotheses. The proof of Proposition 3.6 uses parts of Brendle’s original blowdown argument, decay estimates of Dung–Sung [dung-sung] and Zhao–Zhu [zz], tensor identities on Ricci solitons (where $h$ being a Lie derivative is essential), maximum principles, and standard integration-by-parts arguments.
In §4, we use the results of §3 to establish global symmetry principles. Under mild conditions, we show that the isometry group of each level set of $f$ acts isometrically on the whole of $(M,g)$, and this action preserves all level sets of $f$ (Lemma 4.1). Moreover, if the level sets are homogeneous, then the orbit space is one-dimensional (Proposition 4.2). In this case, standard facts about cohomogeneity one manifolds apply to pin down the global structure of $(M,g,f)$.
In §5, we prove the uniqueness of the Bryant soliton (Theorem 1.4) following [brendle2013rotational, brendle2014rotational, zz]. Given an asymptotically cylindrical steady GRS $(M^{n},g,f)$, we use the results of §3 to find $\frac{n(n-1)}{2}$ Killing fields on each far enough level set of $f$, which implies these level sets are isometric to round spheres. From this, we deduce that $(M,g)$ is globally rotationally symmetric. This last step was not detailed by Brendle and Zhao–Zhu; we give its proof for completeness, and also because a slight generalization of it (Proposition 5.2) is needed. We also prove the non-existence of asymptotically $\mathbb{RP}^{n-1}$-cylindrical steady GRSs (Theorem 1.6). This proceeds almost identically to the proof of Theorem 1.4, except the last step uses Proposition 5.2 to obtain a contradiction.
In §6, we prove the uniqueness of Appleton’s solitons (Theorem 1.8). Let $(M,g,f)$ be an asymptotically $L_{m,k}$-cylindrical steady GRS with $m\geq 1$, $k\geq 3$. The results of §3 induce $(m+1)^{2}$ Killing fields on each far enough level set of $f$. Using the classification of homogeneous metrics on spheres [ziller-homog] and some group theory, we deduce that the level sets are isometric to quotiented Berger spheres $(L_{m,k},g_{a,b})$. The results of §4 imply that $(M,g)$ has a cohomogeneity one isometric action of $U(m+1)$, and is specified outside a submanifold $X$ by Appleton’s ODE system (1.7)–(1.9). The cohomogeneity one formalism classifies the possible diffeomorphism types of $X$; in our case, we find that the metric completes smoothly if and only if $X\cong\mathbb{CP}^{m}$ and the boundary conditions (1.10) are met. We conclude that $(M,g,f)\in\mathcal{A}_{m,k}$.
In §7, we finish the classification of asymptotically $SU(2)/\Gamma$-cylindrical steady GRSs (Theorems 1.10 and 1.11). Here we also describe the solitons $A/\mathbb{Z}_{2}$ (for $A\in\mathcal{A}_{1,2k}$) in more detail. Other than this, the ideas and methods involved are similar to those used in §5 and §6, so we omit an outline for brevity.
Appendix A contains a proof that Appleton’s solitons in $\mathcal{A}_{m,k}$ are asymptotically $L_{m,k}$-cylindrical.
@@ 2. Geometry of asymptotically $\mathcal{S}$-cylindrical steady GRSs
Throughout this section, $(M^{n},g,f)$ denotes an asymptotically $\mathcal{S}$-cylindrical steady GRS of dimension $n\geq 3$, where $\mathcal{S}$ is a spherical space form of dimension $n-1$.
@@ 2.1. Steady soliton identities
(2.5) $\Delta_{f}f=1$
which will be repeatedly used throughout.
@@ 2.2. Geometric properties
By [deng-zhu]*Lemma 6.5, asymptotic $\mathcal{S}$-cylindricity implies that $R=o(1)$ as $r\to\infty$. Then the claim follows from (2.3) and Lemma 2.1. ∎
@@
It is clear that $(M,g)$ does not split off a line. Hence:
Theorem 2.4 ([mw-smms]*Corollary 1.1).
@@ 3. Killing fields
Given an asymptotically $\mathcal{S}$-cylindrical steady GRS $(M^{n},g,f)$, we will show in this section that every symmetry of the asymptotic cross-section $(\mathcal{S},\bar{g}_{\infty})$ manifests as a genuine symmetry of $(M,g)$. More precisely, we prove:
@@
(3.3) $\sup_{\{f=\rho\}}|W_{i}-U_{i}|_{\bar{g}_{\infty}}^{2}$ $\leq\sup_{\{f=\rho\}}|W_{i}-U_{i}|_{\bar{g}_{\infty}-\frac{1}{2(n-2)\rho}g_{\rho}}^{2}+\sup_{\{f=\rho\}}|W_{i}-U_{i}|_{\frac{1}{2(n-2)\rho}g_{\rho}}^{2}$
(3.4) $\leq\mathcal{O}(\rho^{-\frac{1}{2}-\frac{1}{2}\delta})\sup_{\{f=\rho\}}|W_{i}-U_{i}|_{\bar{g}_{\infty}}^{2}+\mathcal{O}(\rho^{-1}).$
Hence,
@@ 3.1. Reduction to a Liouville theorem for symmetric 2-tensors
For the rest of this section, we fix an asymptotically $\mathcal{S}$-cylindrical steady GRS $(M^{n},g,f)$. The goal is to prove Theorem 3.1. First, we pull back Killing fields from $(\mathcal{S},\bar{g}_{\infty})$ to obtain approximate Killing fields on $(M,g)$.
@@
so the function $\mathcal{L}_{\tilde{W}}f=\langle\tilde{W},\nabla f\rangle$ is constant. Since $f$ attains a minimum by Lemma 2.1, we conclude that $\langle\tilde{W},\nabla f\rangle$ vanishes identically. From this, and the identification (2.9) of the end of $(M,g)$, one checks (e.g. by a computation in coordinates) that $\mathcal{L}_{\tilde{W}}g_{\rho}=\mathcal{L}_{\tilde{W}}g=0$ for each large $\rho$. Thus $\tilde{W}$ is a Killing field on $(\mathcal{S},g_{\rho})$. ∎
@@ 4. Global symmetry and cohomogeneity one structure
Next, we establish some global symmetry principles. For a vector field $X$ on a complete Riemannian manifold $(M,g)$, let $\Phi^{X}$ be its time 1 flow. Since $(M,g)$ is complete, $\Phi^{X}$ is a well-defined isometry of $(M,g)$ for any Killing field $X$. Let $\mathcal{I}(M,g)$ be the isometry group of $(M,g)$, and $\mathcal{I}^{0}(M,g)$ its identity component. These are finite-dimensional Lie groups by the Myers–Steenrod theorem.
@@
    The singular orbit $X=\pi^{-1}(0)$ is diffeomorphic to $G/K$, where $K$ is a closed subgroup of $G$ such that $H\subset K$ and $K/H\cong\mathbb{S}^{\ell}$ for some $\ell\geq 0$.
@@
    $M$ is diffeomorphic to the total space of the vector bundle $E=G\times_{K}\mathbb{R}^{\ell+1}$ over $G/K\cong X$, where $K$ acts via $k\cdot(g,v)=(gk^{-1},kv)$, and $kv$ is the radial extension of the action of $K$ on $K/H\cong\mathbb{S}^{\ell}\hookrightarrow\mathbb{R}^{\ell+1}$.
@@
As $\pi:M\to M/G$ is continuous and $M$ is noncompact, we can rule out $\mathbb{S}^{1}$ and $[0,1]$. We can also rule out $(0,1)$, as by [mostert] this would imply $M$ is diffeomorphic to $(0,1)\times\mathcal{S}$, contradicting Theorem 2.4. Hence, $M/G$ is homeomorphic to $[0,1)$.
Thus, we have proved (a) and (b). The other assertions are standard facts about manifolds admitting an isometric action of cohomogeneity one with orbit space homeomorphic to $[0,1)$; see for instance [alexandrino-bettiol]*§6.3 or [hoelscher, mostert]. ∎
5. Uniqueness of the Bryant soliton
5.1. Proof of Theorem 1.4
We now prove the uniqueness of the Bryant soliton, exactly as in [brendle2013rotational, brendle2014rotational, zz].
@@ 5. Uniqueness of the Bryant soliton
Let $(M^{n},g,f)$ be an asymptotically cylindrical steady GRS of dimension $n\geq 3$. Then $(\mathbb{S}^{n-1},\bar{g}_{\infty})$ (where $\bar{g}_{\infty}$ is from Lemma 2.6) has $\frac{n(n-1)}{2}$ linearly independent Killing fields $\{\bar{U}_{i}\}_{i=1}^{n(n-1)/2}$. Thus, Corollary 3.2 gives $\frac{n(n-1)}{2}$ linearly independent Killing fields $\{W_{i}\}_{i=1}^{n(n-1)/2}$ on $\{f=\rho\}$ for all large $\rho$. It follows that these level sets have constant sectional curvature. The conclusion follows from Proposition 5.2 below. ∎
@@ 5.2. Roundness near infinity implies global roundness
where $\{\sigma_{1},\sigma_{2},\sigma_{3}\}$ is the dual coframe of a left-invariant frame $\{X_{1},X_{2},X_{3}\}$ satisfying
@@
If all far enough level sets of $f$ have constant sectional curvature, then $\mathcal{S}=\mathbb{S}^{n-1}$ and $(M,g)$ is the Bryant soliton.
@@
where $\{g(s)\}_{s>0}$ is a family of $G$-invariant metrics on a fixed copy of $G/H\cong\mathcal{S}$. From the description of $(M,g)$ as a vector bundle over $X$, the metric spaces $(\mathcal{S},g(s))$ converge in the Hausdorff sense as $s\to 0$ to $X$ with the metric stucture from $g$. Moreover, $f$ is a function of $s$, and by Lemma 2.1, $f\sim s$ at infinity.
Since all far enough level sets of $f$ have constant sectional curvature, it follows from Lemma 5.1 that
@@
(See for instance [chow2023ricci]*§6, noting the difference in sign convention for $f$.) Our steady GRS $(M,g,f)$ therefore gives a solution $(w,f)$ to this system on $(\rho_{0},\infty)$. Let $(a,\infty)$, $-\infty\leq a<\rho_{0}$, be the maximal interval on which this solution extends while maintaining $w>0$. The metric $\hat{g}=ds^{2}+w(s)^{2}\bar{g}_{\mathcal{S}}$ on $(a,\infty)\times\mathcal{S}$, being a gradient Ricci soliton, is real-analytic [ivey-analytic]. Likewise, $g$ is real-analytic on $(0,\infty)\times\mathcal{S}$. We have $g=\hat{g}$ on $(\rho_{0},\infty)\times\mathcal{S}$. Hence, unique continuation results (e.g. [kobayashi-nomizu-1]*Lemma §VI.6.3’) apply to give
@@
If $a<0$, then $w(s)$ has a positive limit as $s\to 0$, so $(\mathcal{S},g(s))$ converges in the Hausdorff sense as $s\to 0$ to a limiting space homeomorphic to $\mathcal{S}$. This is a contradiction since $X$ is not homeomorphic to $\mathcal{S}$. Thus $a\geq 0$. We further claim that
@@
By (5.7) and (5.8), the metrics $g(s)=w(s)^{2}\bar{g}_{\mathcal{S}}$ degenerate to a point as $s\to a^{+}$. If $a>0$, this gives a contradiction in view of (5.3). Hence $a=0$, and $X$ is a point. In order for $(M,g)$ to be smooth, we require that $\mathcal{S}=\mathbb{S}^{n-1}$. In this case, since $(M,g,f)$ is a complete steady GRS by assumption, and is globally rotationally symmetric, it must be the Bryant soliton. ∎
5.3. Proof of Theorem 1.6
This is almost identical to the proof of Theorem 1.4.

### v2

@@ 1.5. Outline of the paper
In §2, we collect known facts about asymptotically $\mathcal{S}$-cylindrical steady GRSs. In particular, any such soliton $(M,g,f)$ opens up like a paraboloid, and the level sets of $f$ converge smoothly (after a certain rescaling) to a spherical space form $(\mathcal{S},\bar{g}_{\infty})$.
@@
where $\operatorname{Rm}(h)_{ik}=\operatorname{Rm}_{ijkl}h^{jl}$ is the action of the curvature tensor. We establish this for the tensor $h=\mathcal{L}_{W}g$ (Proposition 3.6). In contrast to Brendle and Zhao–Zhu [zz], this holds without any extra curvature hypotheses.
In §4, we use the results of §3 to establish global symmetry principles. We show that the isometry group of each level set of $f$ acts isometrically on the whole of $(M,g)$, and this action preserves $f$ (Lemma 4.1). If the level sets of $f$ are homogeneous, then $(M,g)$ admits a cohomogeneity one structure (Proposition 4.2).
In §5, we prove uniqueness of the Bryant soliton (Theorem 1.4) following [brendle2013rotational, brendle2014rotational, zz]. Given an asymptotically cylindrical steady GRS $(M^{n},g,f)$, we use the results of §3 to find $\frac{n(n-1)}{2}$ Killing fields on each far enough level set of $f$, which implies these level sets are isometric to round spheres. From this, we deduce that $(M,g)$ is globally rotationally symmetric. By similar arguments, we prove the non-existence of asymptotically $\mathbb{RP}^{n-1}$-cylindrical steady GRSs on smooth manifolds (Theorem 1.6), as well as parts (a) and (b) of Corollary 1.10 in the orbifold setting.
In §6, we prove the uniqueness of Appleton’s solitons (Theorem 1.8). Let $(M,g,f)$ be an asymptotically $L_{m,k}$-cylindrical steady GRS with $m\geq 1$, $k\geq 3$. The results of §3 induce $(m+1)^{2}$ Killing fields on each far enough level set of $f$. From this, we deduce that the level sets are isometric to quotiented Berger spheres $(L_{m,k},g_{a,b})$. The results of §4 imply that $(M,g)$ has a cohomogeneity one isometric action of $U(m+1)$, and is specified outside a submanifold $X$ by Appleton’s ODE system (1.6)–(1.8). If $M$ is a smooth manifold, then we deduce that $X\cong\mathbb{CP}^{m}$ and the boundary conditions (1.9) are met. We conclude that $(M,g,f)\in\mathcal{A}_{m,k}$, proving Theorem 1.8. In the 4-dimensional orbifold setting, we show that $X$ could also be a point, in which case $(M,g,f)$ is the Bryant soliton quotiented by $\mathbb{Z}_{k}$. This establishes part (c) of Corollary 1.10.
In §7, using similar arguments as in §5 and §6, we finish the classification of asymptotically $SU(2)/\Gamma$-cylindrical steady GRSs on orbifolds (Theorems 1.11 and 1.12).
@@ 2. Geometry of asymptotically $\mathcal{S}$-cylindrical steady GRSs
In this section, we assume $(M^{n},g,f)$ is an asymptotically $\mathcal{S}$-cylindrical steady GRS of dimension $n\geq 3$, where $\mathcal{S}$ is a spherical space form of dimension $n-1$. Throughout this paper, all asymptotically $\mathcal{S}$-cylindrical GRSs are assumed to have the following regularity unless otherwise stated:
  • If $n\neq 4$, then $M$ is a smooth manifold.
  • If $n=4$, then $M$ is a smooth manifold, or it is a smooth orbifold with isolated singularities and a singularity model in the sense of [bamler-3]*Definition 1.1.
@@ 2.1. Steady soliton identities
(2.5) $\Delta_{f}f=1.$
@@ 2.2. Geometric properties
By [deng-zhu]*Lemma 6.5, asymptotic $\mathcal{S}$-cylindricity implies that $R=o(1)$ as $r\to\infty$. This is also true in the 4-dimensional orbifold case [bamler-3]*Theorem 1.3. Then the claim follows from (2.3) and Lemma 2.1. ∎
@@
Theorem 2.4.
@@
Proof.
For smooth manifolds, this follows from [mw-smms]*Corollary 1.1 since $(M,g)$ clearly does not split off a line. In the 4-dimensional orbifold case, see [bamler-3]*Theorem A.1. ∎
@@ 3. Killing fields
Given an asymptotically $\mathcal{S}$-cylindrical steady GRS $(M^{n},g,f)$, we will show in this section that every continuous symmetry of the asymptotic cross-section $(\mathcal{S},\bar{g}_{\infty})$ manifests as a genuine symmetry of $(M,g)$. More precisely, we prove:
@@ 3.1. Reduction to a Liouville theorem for symmetric 2-tensors
For the rest of this section, we fix an asymptotically $\mathcal{S}$-cylindrical steady GRS $(M^{n},g,f)$. First, we pull back Killing fields from $(\mathcal{S},\bar{g}_{\infty})$ to obtain approximate Killing fields on $(M,g)$.
@@
so the function $\mathcal{L}_{\tilde{W}}f=\langle\tilde{W},\nabla f\rangle$ is constant. Since $f$ attains a minimum by Lemma 2.1, we conclude that $\langle\tilde{W},\nabla f\rangle$ vanishes identically. From this, and the identification (2.9) of the end of $(M,g)$, one checks that $\mathcal{L}_{\tilde{W}}g_{\rho}=\mathcal{L}_{\tilde{W}}g=0$ for each large $\rho$. Thus $\tilde{W}$ is a Killing field on $(\mathcal{S},g_{\rho})$. ∎
@@ 4. Global symmetry and cohomogeneity one structure
Next, we establish some global symmetry principles. For a vector field $X$ on $M$, let $\Phi^{X}$ be its time 1 flow. If $(M,g)$ is complete, then $\Phi^{X}$ is a well-defined isometry of $(M,g)$ for any Killing field $X$. Let $\mathcal{I}(M,g)$ be the isometry group of $(M,g)$, and $\mathcal{I}^{0}(M,g)$ its identity component. These are finite-dimensional Lie groups by the Myers–Steenrod theorem (see [isom-orb]*Theorem 1 for the orbifold version).
@@
    The singular orbit $X=\pi^{-1}(0)$ is diffeomorphic to $G/K$, where $K$ is a closed subgroup of $G$ such that $H\subset K$ and $K/H\cong\mathcal{S}^{\prime}$ for some spherical space form $\mathcal{S}^{\prime}$ of dimension $\ell\geq 0$. Moreover, $M$ is a smooth manifold if and only if $\mathcal{S}^{\prime}$ is a sphere.
@@
    $M$ is diffeomorphic to the orbifiber bundle $E=G\times_{K}\mathrm{Cone}(\mathcal{S}^{\prime})$ over $G/K\cong X$, where $K$ acts via $k\cdot(g,v)=(gk^{-1},kv)$, and $kv$ is the radial extension of the action of $K$ on $K/H\cong\mathcal{S^{\prime}}$.
@@
(This is also true on Riemannian orbifolds, as is implied by the results in [isom-alexandrov]*§2 which apply in this setting.) As $M$ is noncompact, we can rule out $\mathbb{S}^{1}$ and $[0,1]$. We can also rule out $(0,1)$, as by [mostert] this would imply $M$ is diffeomorphic to $(0,1)\times\mathcal{S}$, contradicting Theorem 2.4. Hence, $M/G$ is homeomorphic to $[0,1)$.
Thus, we have proved (a) and (b). The other assertions are well-known facts about manifolds admitting a cohomogeneity one isometric action with orbit space $[0,1)$; see e.g. [alexandrino-bettiol]*§6.3 and [hoelscher]*§1. For orbifolds, we refer to [gonzalez-alvaro-phd]. ∎
5. Uniqueness of the Bryant soliton and its $\mathbb{Z}_{2}$-quotient
5.1. Uniqueness of the Bryant soliton
We proceed exactly as in [brendle2013rotational, brendle2014rotational, zz].
@@ 5.1. Uniqueness of the Bryant soliton
Let $(M^{n},g,f)$ be an asymptotically cylindrical steady GRS on a smooth manifold of dimension $n\geq 3$. Then $(\mathbb{S}^{n-1},\bar{g}_{\infty})$ (where $\bar{g}_{\infty}$ is from Lemma 2.6) has $\frac{n(n-1)}{2}$ linearly independent Killing fields $\{\bar{U}_{i}\}_{i=1}^{n(n-1)/2}$. Thus, Corollary 3.2 gives $\frac{n(n-1)}{2}$ linearly independent Killing fields on $\{f=\rho\}$ for all large $\rho$. It follows that these level sets have constant sectional curvature. The conclusion follows from Proposition 5.2(a) below. ∎
Part (a) of Corollary 1.10 is proved using the same argument, but instead using part (b) of Proposition 5.2 in the last step.
@@ 5.2. Roundness near infinity implies global roundness
where $\{\sigma_{1},\sigma_{2},\sigma_{3}\}$ is dual to a left-invariant frame $\{X_{1},X_{2},X_{3}\}$ satisfying
@@
Suppose that all far enough level sets of $f$ have constant sectional curvature. Then we have the following conclusions depending on the regularity of $M$.
  1. (a)
    If $M$ is a smooth manifold, then $\mathcal{S}=\mathbb{S}^{n-1}$ and $(M,g,f)$ is the Bryant soliton.
  2. (b)
    If $n=4$ and $M$ is a smooth orbifold with isolated singularities, then $(M,g,f)$ is the corresponding quotient of the Bryant soliton.
@@
where $\{g(s)\}_{s>0}$ is a family of $G$-invariant metrics on a fixed copy of $G/H\cong\mathcal{S}$. From the description of $(M,g)$ as a bundle over $X$, the metric spaces $(\mathcal{S},g(s))$ converge in the Hausdorff sense as $s\to 0$ to $X$ with the submanifold metric. Moreover, $f$ is a function of $s$, and by Lemma 2.1, $f\sim s$ at infinity.
Since all far enough level sets of $f$ have constant sectional curvature, Lemma 5.1 gives
@@
(See e.g. [chow2023ricci]*§6, noting the difference in sign convention for $f$.) Our steady GRS $(M,g,f)$ therefore gives a solution $(w,f)$ to this system on $(\rho_{0},\infty)$. Let $(a,\infty)$, $-\infty\leq a<\rho_{0}$, be the maximal interval on which this solution extends while maintaining $w>0$. The metric $\hat{g}=ds^{2}+w(s)^{2}\bar{g}_{\mathcal{S}}$ on $(a,\infty)\times\mathcal{S}$, being a gradient Ricci soliton, is real-analytic [ivey-analytic]. Likewise, $g$ is real-analytic on $(0,\infty)\times\mathcal{S}$. We have $g=\hat{g}$ on $(\rho_{0},\infty)\times\mathcal{S}$. Hence, unique continuation (e.g. [kobayashi-nomizu-1]*Lemma §VI.6.3’) applies to give
@@
If $a<0$, then $w(s)$ has a positive limit as $s\to 0$, so $(\mathcal{S},g(s))$ converges in the Hausdorff sense as $s\to 0$ to a space homeomorphic to $\mathcal{S}$. This is a contradiction since $X$ is not homeomorphic to $\mathcal{S}$. Thus $a\geq 0$. We further claim that
@@
By (5.7) and (5.8), the metrics $g(s)=w(s)^{2}\bar{g}_{\mathcal{S}}$ degenerate to a point as $s\to a^{+}$. Hence $a=0$, and $X$ is a point. By Proposition 4.2(d), $M$ is diffeomorphic to the cone over $\mathcal{S}$. To conclude, we divide into two cases:
  1. (a)
    If $M$ is a smooth manifold, we must have $\mathcal{S}=\mathbb{S}^{n-1}$. In this case, since $(M,g,f)$ is a complete steady GRS by assumption, and is globally rotationally symmetric, it is the Bryant soliton.
  2. (b)
    If $M$ is a 4-dimensional smooth orbifold with isolated singularities, then its universal orbifold cover $\tilde{M}$ is a smooth manifold diffeomorphic to $\mathbb{R}^{4}$. Lifting the soliton structure gives an asymptotically cylindrical steady GRS $(\tilde{M},\tilde{g},\tilde{f})$. By Theorem 1.4, $(\tilde{M},\tilde{g},\tilde{f})$ is the Bryant soliton.
5.3. Uniqueness of the $\mathbb{Z}_{2}$-quotient of the Bryant soliton
The next proof is almost identical to the proof of Theorem 1.4.